OpenStudy (anonymous):

For every real number x (regardless of whether x is positive, negative, or zero), sqrt(x^2)=x.

OpenStudy (anonymous):

sqrt(x^2) = x^(2/2)=x

OpenStudy (anonymous):

But it doesn't work for x being negative, so it's false :)